OpenStudy (anonymous):

A system of equations is shown below: -3x + 7y = -16 -9x + 5y = 16 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. (6 points) Part B: Show that the equivalent system has the same solution as the original system of equations. (4 points)

OpenStudy (anonymous):

@MentionedLyric @Heartstar27 @triangleguru @undeadknight26 i need just theanswers please

OpenStudy (anonymous):

@Masonatttor @KlOwNlOvE @panda_lover @Dobby1 i just want the answer please!

OpenStudy (anonymous):

@PhotoFreak23 @TSwizzle @littlekitty16 @ARMYRANGERS @eric_d @CloverRacer I just want the answers please

OpenStudy (anonymous):

@yayanieves @Heartstar27 @kirbykirby @Catseyeglint911 @Ashy98 @misz.lopez115 @zBrandz23 @Logan319 PLEASE HELP! i just want the answers

OpenStudy (anonymous):

@Haseeb96 @tracy0793 @inkyvoyd @ponchinrin @KellyRJr @zephy @fkdhfdakjfve PLEAZE HELP! i just need the answers i have been on this for over 2h

OpenStudy (anonymous):

Part A: I replace 2nd equation with sum of 1st and 2nd equation so I get system: -3x + 7y = -16 -12x + 12y = 0

OpenStudy (anonymous):

Part B - You must solve original system of equation; you will get solution: x = -4 y = -4 And then substitute this solution into equivalent system of equation: -3*(-4) + 7*(-4) = -16 -12*(-4) + 12*(-4) = 0 And evaluate expressions to prove that equivalent system of equation have same solution.

OpenStudy (anonymous):

OMFG TY TY TY TY

OpenStudy (anonymous):

is this the final answer for B? @triangleguru ?

OpenStudy (anonymous):

B - I think thats final solution. In A part I constructed replaced equation sa B = 1*A+B. So multiple coeficient was "1". If you choose B = 2*A+B, you will get equivalent system: -3x + 7y = -16 -15x + 19y = -16

OpenStudy (anonymous):

okay TY SO VERY MUCH!!! no one else was helping me i been on this question 2h plus